# Multi-Objective Recommendation via Multivariate Policy Learning

### v1

@@ Abstract.
journalyear: 2024copyright: none
@@ 1. Introduction & Motivation
Another recent tendency is that such systems are increasingly often framed as decision-making instead of prediction systems. Indeed, predictions about user-item affinities typically serve to inform real-time decisions about recommendations, rankings, advertisements, and so forth. The decision-making lens allows us to reason about consequences of such systems, and has been gaining popularity in recent years (Jeunen et al., 2022). In the context of real-world platforms, it allows us to frame changes in key online metrics as the consequences of recommendation decisions, which in turn allows us to optimise those online metrics directly (Jeunen, 2021). Approaches that leverage the decision-making literature and frame recommendation as a bandit learning task have led to several practical successes (McInerney et al., 2018; Chen et al., 2019, 2021, 2022; Dong et al., 2020; Ma et al., 2020; Mehrotra et al., 2020; Bendada et al., 2020; Su et al., 2024; Briand et al., 2024). Most common is the off-policy or counterfactual family of approaches, as they allow practitioners to learn and evaluate models offline before online deployment (van den Akker et al., 2024). Such methods typically leverage some form of algorithmic pessimism, optimising a lower bound on the reward they will yield (Jeunen and Goethals, 2021a, 2023; Swaminathan and Joachims, 2015a).
@@
In this work, we propose a general multi-objective recommendation approach that frames the optimisation of the weights that are given to individual objectives as a decision-making problem. To this end, we leverage the Counterfactual Risk Minimisation (CRM) principle to optimise a lower bound on the reward a recommendation policy will yield, thereby extending it to a multivariate continuous action domain (Swaminathan and Joachims, 2015a). We show that existing approaches to construct the lower bound, based on the Central Limit Theorem, provide insufficient coverage in finite sample scenarios, and propose an efficient and effective policy-dependent correction for this problem. In doing so, we combine existing elements in the literature to derive a novel estimator for the Effective Sample Size (Martino et al., 2017; Owen, 2013).
@@ 2. Multi-Objective Recommendation
Users interact with content on the web through various modalities. Online content marketplaces such as Instagram, Reddit, TikTok or ShareChat allow users to view, click, (dis)like, save, share and comment on items that are presented to them by the recommender system. Such systems are typically not optimised for a single type of interaction, but rather aim to maximise a set of positive behaviours through multi-objective optimisation techniques. Decidedly the most common approach in practice is scalarisation (Hwang and Masud, 1979), where the multi-objective optimisation problem is recast with a single scalar objective. Different parameterisations for the function that maps the multiple objectives to a single scalar, then give rise to different Pareto-optimal solutions. When the Pareto-front is convex, a linear scalarisation function is sufficient to represent all Pareto-optimal solutions. Naturally, this does not tell us how the weights of the linear function should be chosen, or where on the Pareto-front the business should position itself.
Formally, assume we need to decide on a recommendation $i\in\mathcal{I}$ to show to a user $u\in\mathcal{U}$, in an attempt to optimise $d\in\mathbb{N}$ objectives represented as $f(u,i)$. A multi-objective decision-making procedure with linear scalarisation can then be described as finding:
@@
Naturally, in top-$n$ settings we can replace the $\max$-operation with a ${\tt sort}$-operation. This type of procedure is exceedingly common in deployed recommender systems on the web. Indeed: public accounts describe Facebook and TikTok’s systems in this way (Smith, 2021; Merrill and Oremus, 2021) and Twitter even open-sourced their exact weights (Twitter, 2023). Mehrotra et al. (Mehrotra et al., 2020) describe how Spotify leverages the Generalised Gini Index (Busa-Fekete et al., 2017) aggregation function to decide on the weights, motivated as preserving fairness between user- and artist-centric objectives. Zhang et al. describe a reinforcement learning approach for Tencent that aims to find weights that maximise long-term user satisfaction (Zhang et al., 2022). Milli et al. study strategic behaviour of users and content providers that might result from the weights being chosen Milli et al. (2023). Jannach and Abdollahpouri present a taxonomy of types of objectives that might be considered in multi-objective recommendation settings, whilst outlining open challenges (Jannach and Abdollahpouri, 2023).
@@ 4. Improving the CRM Lower Bound
We require accurate variance estimates for two reasons: (1) during evaluation, we want to ensure statistical significance of our results, obtaining confidence intervals that exhibit the expected coverage levels so we can make meaningful statements about Type-I errors, and (2) during learning, we make use of sample variance penalisation schemes that assume accurate variance estimates to provide sensible lower bounds. As such: we require an estimator for ${\rm Var}\left(\widehat{V}_{\rm(SN)IPS}(\pi_{\theta},\mathcal{D})\right)$ that implies confidence intervals with expected coverage. That is, when $\lambda\approx 1.96$ in Eq. 7, the lower bound holds in 95% of cases. Traditionally used variance estimators are based on Gaussian approximations, motivated through the CLT. For $\widehat{V}_{\rm IPS}$, the sample variance with $N$ as the sample size is given by:
@@ 4.1. Effective Sample Size Corrections
From this formula, it is clear that $\widetilde{N}$ is monotonically increasing as a function of $N$, and that $\widetilde{N}\equiv N$ when $\widehat{\rm ESS}\equiv N$. In Eqs. 910, we then simply replace $N$ with $\widetilde{N}$ and denote this as $\widehat{{\rm Var_{ESS}}}$. This gives rise to a symmetric confidence interval as:
@@ 5.1. Violating the common support assumption
The main problem that this violation implies, is that the expected importance weight no longer equals 1 (London and Joachims, 2022). Aside from the fact that this implies that the conventional IPS estimator is now biased, it brings along additional problems. Indeed, baseline corrections now have a significant effect on the estimator, and should be avoided. Usually, practitioners recenter observed rewards as a simple variance-reduction technique. When the common support assumption holds, we can straightforwardly show that any constant scalar translation $\beta$ preserves the unbiasedness of the estimator:
@@ 5.2. Self-normalisation becomes unstable
The SNIPS estimator introduced in Eq. 5 also leverages the fact that $\mathbb{E}_{a\sim\pi_{0}}\left[\frac{\pi_{\theta}(a)}{\pi_{0}(a)}\right]=1.$ Because this equivalence no longer holds, the SNIPS estimator is no longer unbiased. More problematically, when leveraging this estimator for learning purposes, we are incentivised to find a policy that minimises $\mathbb{E}_{a\sim\pi_{0}}\left[\frac{\pi_{\theta}(a)}{\pi_{0}(a)}\right]$ (as it appears in the denominator of Eqs. 5 and 12). This quantity can become arbitrarily small by moving $\pi_{\theta}$ to or beyond the borders of the logging policy’s domain. Indeed: this decreases $\mathbb{E}_{a\sim\pi_{0}}\left[\frac{\pi_{\theta}(a)}{\pi_{0}(a)}\right]$ and increases $\widehat{V}_{\rm SNIPS}$ as a result. It should nevertheless intuitively be clear that this is undesirable behaviour: placing probability density away from the observed data should not increase our reward estimate. This is closely related to the “propensity overfitting” phenomenon described by Swaminathan and Joachims (Swaminathan and Joachims, 2015b), which the uniform logging policy cannot avoid here. As such, we cannot enjoy the variance reduction that self-normalisation brings without introducing a statistical bias that is problematic.
@@ 5.3. The curse of dimensionality
Intuitively, what practitioners expect when choosing a uniformly distributed data collection policy is that he probability of a point falling into a region $[\mu-\epsilon,\mu+\epsilon]$ (where $\mu$ is the centre of the distribution) grows linearly with $\epsilon$ (until the upper bound of the interval). This is clearly valid for the univariate case. Nevertheless, this is not what we get for the multivariate case. Indeed, this probability can be computed as $\mathsf{P}(x\in[\mu-\epsilon,\mu+\epsilon])=(1-2\epsilon)^{d}$, implying that a hypersphere with a radius covering 80% of the maximal distance to the origin ($\epsilon=0.4$), only covers roughly 16% of the data points for $d=8$. This fraction decreases monotonically as $d$ grows.
@@
Furthermore, common support for the logging and target policies opens the door to many statistical tools and tricks that allow for more effective and efficient estimation for both evaluation and learning: baseline corrections, and self-normalisation.
@@ 6. Designing a Sensitive Reward signal
In essence, we wish to use a reward signal that is highly (causally) correlated with the North Star, but potentially has lower variance. Variance reduction techniques are a staple in the online experimentation research literature, typically leveraging regression adjustments as additive control variates (Deng et al., 2013; Poyarkov et al., 2016; Budylin et al., 2018; Baweja et al., 2024). Note that this line of work is analogous to the use of Doubly Robust estimators in the off-policy learning literature (Dudík et al., 2014). As these techniques would require us to introduce an additional model that estimates the reward, they are out-of-scope for the purposes of this work.
@@ 7.2. Improved Recommendations (RQ2–4)
The base recommendation model on the platforms is a Multi-gate Mixture of Experts (MMOE) multi-task learning model (Ma et al., 2018), used to generate predictions for 9 different behaviour signals (e.g. like, share, view, $\ldots$). These are then combined following Eq. 1.
@@ 7.2.2. Offline Results (RQ2–3)
To obtain an offline dataset to learn from, we randomise weights for 4.5 million users and log their behaviour on the platform over a period of three weeks. The randomised weights correspond to the actions, and the behaviours inform the rewards we wish to optimise using our proposed approach.
We optimise a deterministic policy to maximise a lower bound on the $\widehat{V}_{\rm SNIPS}$ estimator, given by Eq. 12. For the learnt policy, we visualise the obtained ESS-corrected confidence interval for the value estimates when we vary the bandwidth of the Gaussian kernel (i.e. the standard deviation $\sigma$). These results are visualised in Figure 4. The policy we wish to deploy is deterministic, i.e. $\sigma\equiv 0$. Naturally, this would lead to an extremely low ESS, and high variance as a result. As we increase $\sigma$, we observe that the variance of the estimator decreases as its bias increases. Indeed, for large values of $\sigma$, the estimates coincide with the empirical reward obtained by the logging policy, as the widening Gaussian kernel tend to smooth over all observed samples. This visualises the bias–variance trade-off that occurs for off-policy evaluation in continuous action domains.
@@ 7.2.3. Online Results (RQ4)
To reproduce, validate and extend these empirical insights, we perform a second online A/B-test on another large-scale short-video platform, considering similar metrics. In this experiment, we learn different scalarisation weights for different short-video feeds users can interact with on the platform (i.e. home page feed, specific feeds, et cetera). Reassuringly, we also observe that the learnt policy is able to improve the target metric significantly. The time users spend on the platform and user retention also improve significantly. Because these metrics are not easily moved by online experiments, this represents an important improvement to end-users’ experience on the platforms, and signifies the effectiveness of our method.

### v2

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: 18th ACM Conference on Recommender Systems; October 14–18, 2024; Bari, Italybooktitle: 18th ACM Conference on Recommender Systems (RecSys ’24), October 14–18, 2024, Bari, Italydoi: 10.1145/3640457.3688132isbn: 979-8-4007-0505-2/24/10
@@ 1. Introduction & Motivation
Another recent tendency is that such systems are increasingly often framed as decision-making instead of prediction systems. Indeed, predictions about user-item affinities typically serve to inform real-time decisions about recommendations, rankings, advertisements, and so forth. The decision-making lens allows us to reason about consequences of such systems, and has been gaining popularity in recent years (Jeunen et al., 2022). In the context of real-world platforms, it allows us to frame changes in key online metrics as the consequences of recommendation decisions, which in turn allows us to optimise those online metrics directly (Jeunen, 2021; Jeunen et al., 2024). Approaches that leverage the decision-making literature and frame recommendation as a bandit learning task have led to several practical successes (McInerney et al., 2018; Chen et al., 2019, 2021, 2022; Dong et al., 2020; Ma et al., 2020; Mehrotra et al., 2020; Bendada et al., 2020; Su et al., 2024; Briand et al., 2024). Most common is the off-policy or counterfactual family of approaches, as they allow practitioners to learn and evaluate models offline before online deployment (van den Akker et al., 2024). Such methods typically leverage some form of algorithmic pessimism, optimising a lower bound on the reward they will yield (Jeunen and Goethals, 2021a, 2023; Swaminathan and Joachims, 2015a).
@@
In this work, we propose a general multi-objective recommendation approach that frames the optimisation of the weights that are given to individual objectives as a decision-making problem. To this end, we leverage the Counterfactual Risk Minimisation (CRM) principle to optimise a lower bound on the reward a recommendation policy will yield, thereby extending it to a multivariate continuous action domain (Swaminathan and Joachims, 2015a). We show that existing approaches to construct the lower bound, based on the Central Limit Theorem (CLT), provide insufficient coverage in finite sample scenarios, and propose an efficient and effective policy-dependent correction for this problem. In doing so, we combine existing elements in the literature to derive a novel estimator for the Effective Sample Size (Martino et al., 2017; Owen, 2013).
@@ 2. Multi-Objective Recommendation
Users interact with content on the web through various modalities. Online content marketplaces such as Instagram, Reddit, TikTok or ShareChat allow users to view, click, (dis)like, save, share and comment on items that are presented to them by the recommender system. Such systems are typically not optimised for a single type of interaction, but rather aim to maximise a set of positive behaviours through multi-objective optimisation techniques. Decidedly the most common approach in practice is scalarisation (Hwang and Masud, 1979), where the multi-objective optimisation problem is recast with a single scalar objective. Different parameterisations for the function that maps the multiple objectives to a single scalar, then give rise to different Pareto-optimal solutions. Pareto-optimality in multi-objective recommendation applications is a well-studied topic (Ribeiro et al., 2012; Lin et al., 2019; Xie et al., 2021). When the Pareto-front is convex, a linear scalarisation function is sufficient to represent all Pareto-optimal solutions. Naturally, this does not tell us how the weights of the linear function should be chosen, or where on the Pareto-front the business should position itself to optimise long-term business goals.
Formally, assume we need to decide on a recommendation $i\in\mathcal{I}$ to show to a user $u\in\mathcal{U}$, in an attempt to optimise $d\in\mathbb{N}$ objectives represented as $f(u,i)$. A multi-objective decision-making procedure with linear scalarisation weights $a\in\mathbb{R}^{d}$ can then be described as finding the items that satisfy:
@@
Naturally, in top-$n$ settings we can replace the $\max$-operation with a ${\tt sort}$-operation. This type of procedure is exceedingly common in deployed recommender systems on the web. Indeed: public accounts describe Facebook and TikTok’s systems in this way (Smith, 2021; Merrill and Oremus, 2021) and Twitter even open-sourced their exact weights (Twitter, 2023). Mehrotra et al. (Mehrotra et al., 2020) describe how Spotify leverages the Generalised Gini Index (Busa-Fekete et al., 2017) aggregation function to decide on the weights, motivated as preserving fairness between user- and artist-centric objectives. Zhang et al. describe a reinforcement learning approach for Tencent that aims to find weights that maximise long-term user satisfaction (Zhang et al., 2022). Milli et al. study strategic behaviour of users and content providers that might result from the weights being chosen (Milli et al., 2023). Jannach and Abdollahpouri present a taxonomy of types of objectives that might be considered in multi-objective recommendation settings, whilst outlining open challenges (Jannach and Abdollahpouri, 2023).
@@ 4. Improving the CRM Lower Bound
We require accurate variance estimates for two reasons: (1) during evaluation, we want to ensure statistical significance of our results, obtaining confidence intervals that exhibit the expected coverage levels so we can make meaningful statements about probabilities, and (2) during learning, we make use of sample variance penalisation schemes that assume accurate variance estimates to provide sensible lower bounds. As such: we require an estimator for ${\rm Var}\left(\widehat{V}_{\rm(SN)IPS}(\pi_{\theta},\mathcal{D})\right)$ that implies confidence intervals with expected coverage. That is, when $\lambda\approx 1.96$ in Eq. 7, the lower bound holds in 95% of cases. Traditionally used variance estimators are based on Gaussian approximations, motivated through the CLT. For $\widehat{V}_{\rm IPS}$, the sample variance with $N$ as the sample size is given by:
@@ 4.1. Effective Sample Size Corrections
From this formula, it is clear that $\widetilde{N}$ is monotonically increasing as a function of $N$, and that $\widetilde{N}\equiv N$ when $\widehat{\rm ESS}\equiv N$. Furthermore, when ${\rm ESS}>1$, as $N$ grows, Eq. 11 converges towards $N$. As a result, $\widetilde{N}$ is a consistent estimator for $N$, and as a consequence, variance estimates using $\widetilde{N}$ instead of $N$ are also consistent.
In Eqs. 910, we then simply replace $N$ with $\widetilde{N}$ and denote this as $\widehat{{\rm Var_{ESS}}}$. This gives rise to a symmetric confidence interval as:
@@ 5.1. Violating the common support assumption
The main problem that this violation implies, is that the expected importance weight no longer equals 1 (London and Joachims, 2022). Aside from the fact that this implies that the conventional IPS estimator is now biased, it brings along additional problems. Indeed, baseline corrections now have a significant effect on the estimator, and should be avoided. Usually, practitioners recentre observed rewards as a simple variance-reduction technique (Gupta et al., 2024). When the common support assumption holds, we can straightforwardly show that any constant scalar translation $\beta$ preserves the unbiasedness of the estimator:
@@ 5.2. Self-normalisation becomes unstable
The SNIPS estimator introduced in Eq. 5 also leverages the fact that $\mathbb{E}_{a\sim\pi_{0}}\left[\frac{\pi_{\theta}(a)}{\pi_{0}(a)}\right]=1.$ Because this equivalence no longer holds, the SNIPS estimator is no longer asymptotically unbiased. More problematically, when leveraging this estimator for learning purposes, we are incentivised to find a policy that minimises $\mathbb{E}_{a\sim\pi_{0}}\left[\frac{\pi_{\theta}(a)}{\pi_{0}(a)}\right]$ (as it appears in the denominator of Eqs. 5 and 12). This quantity can become arbitrarily small by moving $\pi_{\theta}$ to or beyond the borders of the logging policy’s domain. Indeed: this decreases $\mathbb{E}_{a\sim\pi_{0}}\left[\frac{\pi_{\theta}(a)}{\pi_{0}(a)}\right]$ and increases $\widehat{V}_{\rm SNIPS}$ as a result. It should nevertheless intuitively be clear that this is undesirable behaviour: placing probability density away from the observed data should not increase our reward estimate. This is closely related to the “propensity overfitting” phenomenon described by Swaminathan and Joachims (Swaminathan and Joachims, 2015b), which the uniform logging policy cannot avoid here. As such, we cannot enjoy the variance reduction that self-normalisation brings without introducing a statistical bias that is problematic.
@@ 5.3. The curse of dimensionality
Intuitively, what practitioners expect when choosing a uniformly distributed data collection policy is that the probability of a point falling into a region $[\mu-\epsilon,\mu+\epsilon]$ (where $\mu$ is the centre of the distribution) grows linearly with $\epsilon$ (until the upper bound of the interval). This is clearly valid for the univariate case. Nevertheless, this is not what we get for the multivariate case. Indeed, this probability can be computed as $\mathsf{P}(x\in[\mu-\epsilon,\mu+\epsilon])=(1-2\epsilon)^{d}$, implying that a hypersphere with a radius covering 80% of the maximal distance to the origin ($\epsilon=0.4$), only covers roughly 16% of the data points for $d=8$. This fraction decreases monotonically as $d$ grows.
@@
Furthermore, common support for the logging and target policies opens the door to many statistical tools and tricks that allow for more effective and efficient estimation for both evaluation and learning: baseline corrections (Gupta et al., 2024), and self-normalisation (Swaminathan and Joachims, 2015b).
@@ 6. Designing a sensitive reward signal
In essence, we wish to use a reward signal that is highly (causally) correlated with the North Star, but potentially has lower variance. Variance reduction techniques are a staple in the online experimentation research literature, typically leveraging regression adjustments as additive control variates (Deng et al., 2013; Poyarkov et al., 2016; Budylin et al., 2018; Baweja et al., 2024). Note that this line of work is analogous to the use of Doubly Robust estimators in the off-policy learning literature (Dudík et al., 2014; Jeunen and Goethals, 2020). As these techniques would require us to introduce an additional model that estimates the reward, they are out-of-scope for the purposes of this work.
@@ 7. Experiments & Discussion
All source code to reproduce the empirical results on synthetic data can be found at github.com/olivierjeunen/multivariate-policy-learning-recsys-2024.
@@ 7.2. Improved Recommendations (RQ2–4)
The base recommendation model on the platforms is a Multi-gate Mixture of Experts (MMOE) multi-task learning model (Ma et al., 2018), used to generate predictions for 9 different behaviour signals (i.e. like, share, favourite, end session and view, among others). These are then combined following Eq. 1.
@@ 7.2.2. Offline Results (RQ2–3)
To obtain an offline dataset to learn from, we randomise weights for 4.5 million users and log their behaviour on the platform over a period of three weeks. The randomised weights correspond to the actions, and the behaviours inform the rewards we wish to optimise using our proposed multivariate off-policy learning approach.
We optimise a deterministic policy to maximise a lower bound on the $\widehat{V}_{\rm SNIPS}$ estimator, given by Eq. 12. For the learnt policy, we visualise the obtained ESS-corrected confidence interval for the value estimates when we vary the bandwidth of the Gaussian kernel (i.e. the standard deviation $\sigma$). These results are visualised in Figure 4. The policy we wish to deploy is deterministic, i.e. $\sigma\equiv 0$. Naturally, this would lead to an extremely low ESS, and high variance as a result. As we increase $\sigma$, we observe that the variance of the estimator decreases as its bias increases. Indeed, for large values of $\sigma$, the estimates coincide with the empirical reward obtained by the logging policy, as the widening Gaussian kernel tends to smooth over all observed samples. This visualises the bias–variance trade-off that occurs for off-policy evaluation in continuous action domains.
@@ 7.2.3. Online Results (RQ4)
To reproduce, validate and extend these empirical insights, we perform a second online A/B-test on another large-scale short-video platform, considering similar metrics. We report results in Table 1(b). In this experiment, we learn different scalarisation weights for different short-video feeds users can interact with on the platform (i.e. home page feed, specific feeds, et cetera). Reassuringly, we also observe that the learnt policy is able to improve the target metric significantly. The time users spend on the platform and user retention also improve significantly. Because these metrics are not easily moved by online experiments, this represents an important improvement to end-users’ experience on the platforms, and signifies the effectiveness of our method.